Given cot x is undefined and cos>0, find csc x
Answer by stanbon(75887)

Given cot x is undefined and cos>0, find csc x
------------------------
cot(x) = x/y
---
If cot(x) is undefined, y = 0
-----------------------------
csc(x) = r/y
So, csc(x) = r/0 is undefined
-------------------------------------
Cheers,
Stan H.